Grade 12th PassPhysical Chemistry

A Gaseous compound of nitrogen and hydrogen contains 12.5% ( by mass) of hydrogen. The density of the compound relative to hydrogen is 16. The molecular formula of the compound is : 1. NH2 2. N3H 3. NH3 4. N2H4

Vignesh M
12 Years agoGrade 12th Pass

2 Answers

Aarti Gupta
ApprovedApproved Tutor Answer12 Years ago
The empirical formula for the compound comes out to be NH2.
Thus empirical formula mass = 14 + 2* 1 = 16
Calculation of molecular formula
n = Molecular mass/Empirical formula mass
Now,
molecular mass = 2 * density
= 2* 16 = 32
Therefore,
n = 32/16 = 2
Thus molecular formula will be = Empirical formula * n
molecular formula = NH2 * 2 = N2H4
ANS. N2H4
Thanks& Regards
Aarti Gupta
askiitians Faculty
Meghana
7 Years ago
N:H=87.5:12.5=1:2Therefore empirical formula is *NH2* E. F weight is 16 gM. F = (MF. Wt÷EF wt)* E. FM. F=(32÷16)*NH2=